# Differentiation Through Black-Box Quadratic Programming Solvers

### v3

@@ 3.3 Extracting derivatives from a QP solver’s solution
Figure 2: Schematic active set differentiation. Left: a QP is shown by its quadratic level sets and polyhedral feasible set; the solution lies on a facet of the boundary; perturbations of the constraints lead to perturbations in the solution. Right: the perturbation of the solution remains the same when inactive constraints are eliminated.
@@
Explicit Active Set Differentiation.
@@
A proof of this Theorem, based on the Basic Sensitivity Theorem [Fiacco, 1976], is provided in Appendix B, along with a calculation of the derivatives using differential matrix calculus [Petersen & Pedersen, 2008; Magnus & Neudecker, 1988]. We note that this result is closely related to analyses studied in multi-parametric programming [Bemporad et al., 2002; Pistikopoulos et al., 2020; Spjøtvold et al., 2006; Arnström & Axehill, 2024; Narciso et al., 2022].
@@
Notably, for quadratic programming, the Basic Sensitivity Theorem allows us to bypass implicit differentiation techniques [Krantz & Parks, 2012]. We emphasize that this does not imply that the general solution or its active set admit a closed-form expression. Rather, while we perform explicit differentiation, the implicit function theorem remains key in establishing the local equivalence between the original and reduced problems -— the resulting derivatives are the same, though the derivations differ. The derivatives in (7) are obtained via ordinary (explicit) differentiation of the closed-form solution to the reduced QP (6), whereas those in (4) arise from implicit differentiation of the full nonlinear KKT conditions (2), followed by pruning inactive constraints. This distinction, formalized in Theorem 3.1, underscores a critical computational insight: once a black-box solver returns the primal solution, the active set can be identified and the derivatives computed directly via (7). Moreover, if the solver returns only the primal solution, the corresponding dual variables can be recovered using (6). Since both (6) and (7) rely on the same KKT matrix $K_{J}$, a single matrix factorization (e.g., via SuperLU [Li, 2005]) suffices for both steps, adding negligible overhead. These insights culminate in the core algorithm of dQP, summarized in Algorithm 1.
@@
Our approach enables compact and efficient gradient computation. The linear system in (7), used to compute both derivatives and dual variables, is symmetric and of reduced size. In contrast, implicit differentiation of the full KKT conditions (1) yields a significantly larger, asymmetric system (3). Beyond simplifying the derivative computation, our approach enables the use of fast, specialized linear solvers that exploit the reduced systems symmetric indefinite KKT matrix structure (e.g., using an LDL factorization as in QDLDL [Stellato et al., 2020; Davis, 2005]).
Figure 3: (a) Schematic of active-set parameter space, coloring regions where the active set is constant. (b) Condition number of the full and reduced linear systems along a line in parameter space.
Empirically, we observe that the reduced linear system (7) is often significantly better conditioned than its full counterpart. Figure 3 illustrates this with an example of a QP governed by two parameters $\theta=(\theta_{1},\theta_{2})$ of [Spjøtvold et al., 2006], calculated using DAQP [Arnström et al., 2022]. Eliminating inactive constraints improves conditioning significantly.
Figure 3 also highlights a key challenge: computing derivatives near non-differentiable singularities where the active set changes and some constraints become weakly active. In such regions, implicit differentiation becomes severely ill-conditioned. Our method is also affected, primarily through difficulty in determining the correct active set at an approximate solution. Several strategies have been proposed for improving active set identification [Cartis & Yan, 2016; Oberlin & Wright, 2006; Burke & Moré, 1988]. Our implementation includes an optional refinement heuristic (Appendix 6), though in all our experiments (Section 4), simple hard thresholding of the primal residual $r_{j}=(Cz^{*}-d)_{j}\geq-\epsilon_{J}$ proved robust.
@@
Our open-source implementation will be publicly released. We implement dQP (Algorithm 1) as a fully differentiable PyTorch module [Paszke et al., 2019], providing an intuitive interface for integrating differentiable QPs into machine learning and bi-level optimization workflows. The implementation supports both dense and sparse problems end-to-end, with appropriate QP and linear solver support. As a PyTorch module, (7) must be rendered as a backward pass; this is described in Appendix C. To ensure modularity, the forward pass supports any QP solver interfaced via the open-source qpsolvers library [Caron et al., 2024b], which provides lightweight access to over 15 commercial and free open-source solvers and easily supports the integration of new ones. We likewise offer flexibility in the choice of linear solver for differentiation, including support for large-scale sparse solvers like Pardiso [Schenk & Gärtner, 2004] and symmetric indefinite solvers like QDLDL [Stellato et al., 2020; Davis, 2005]. For users unsure of which solver to use, dQP includes a profiling tool to help identify the best-performing QP solver for a given problem. Additional implementation details such as constraint normalization, handling non-differentiability, and warm-starting options for bi-level optimization are discussed in Appendix D.
Table 1: Performance of differentiable QP methods for 129 sparse problems from the Maros-Meszaros (MM) dataset [Maros & Mészáros, 1999].
@@ 4 Experimental Results
We have extensively tested dQP to ensure its robustness and evaluate its performance against existing methods for differentiable quadratic programming. Notably, we demonstrate dQP’s strengths in handling large-scale structured and sparse problems, thus complementing custom differentiable GPU-batched solvers such as OptNet [Amos & Kolter, 2017a], which are optimized for solving many small, dense problems simultaneously. Given this focus, and considering the limited availability of state-of-the-art GPU-batchable QP solvers, we conduct our experiments on CPUs, similar to prior works [Bambade et al., 2024; Butler, 2023; Sun et al., 2022; Agrawal et al., 2019a]. Our evaluation includes a large benchmark of over 2,000 challenging sparse and dense QPs taken from public and randomly generated problem datasets, designed to test dQPs robustness and performance. Additional experimental details are provided in Appendix E.
@@
We tested dQP on the QP Benchmark suite [Caron et al., 2024a], focusing on 129 large-scale sparse problems from the standard Maros-Meszaros (MM) dataset [Maros & Mészáros, 1999], which are widely used as stress tests for QP solvers. We compared dQP against other differentiable QP methods available in the PyTorch framework: OptNet [Amos & Kolter, 2017a], QPLayer [Bambade et al., 2024], SCQPTH [Butler, 2023], and CVXPYLayers [Agrawal et al., 2019a]; using the authors’ open-source implementations, each paired with its respective solver. We did not include other general frameworks that either lack direct support for generic QPs in PyTorch, showed subpar performance in our preliminary tests (consistent with findings reported in prior work [Bambade et al., 2024]), or rely on forward passes built on diffcp/CVXPYLayers (see Table 4). For dQP’s forward pass, we selected the top-performing QP solver for each problem, as identified by QP Benchmark [Caron et al., 2024a], to demonstrate the importance of enabling differentiation for the solver that is best suited to each task.
Figure 4: Accuracy versus total forward/backward solve for the Maros-Meszaros dataset [Maros & Mészáros, 1999]. Each point represents a solved problem; point size illustrates dimension; problems solved solely by dQP are circled. The legend shows percentages of success rates; the solvers dQP used and their counts are PIQP 100, Gurobi 9, ProxQP 9, Clarabel 7, OSQP 2, MOSEK 1, QPALM 1.
The scatter plot in Figure 4 shows total runtime (forward + backward), duality gap (accuracy), and problem dimension (illustrated by point size) for each differentiable solver and problem. Aggregate performance across the full dataset and the subset of problems solvable by all methods is summarized in Table 1. CVXPYLayers failed on all but 17 small-scale problems and is therefore excluded from the reported statistics. The MM dataset proved especially challenging, with OptNet and SCQPTH solving fewer than 50% of the problems. In contrast, dQP solved all MM problems and was the only method to succeed on 38 of them (circled in the figure). Moreover, dQP achieved the best runtime and accuracy in 80% and 78% of all problems, respectively. It performed particularly well on large-scale instances (dimension over 1000), being fastest and most accurate in 97% and 93% of such cases. Additional experiments on 450 random dense QPs and 625 sparse QPs (dimensions $10$ to $10^{4}$) are provided in Appendix E.1.3, showing similarly strong performance.
@@
Results for $P_{1}$ are shown in Figure 1, demonstrating dQP’s scalability compared to OptNet and QPLayer. Other methods fail on all but small problems (see Appendix E.1.1). In dimensions greater than 2000, dQP outperforms competing methods by 2-3 orders of magnitude in both speed and accuracy. Competing methods are limited to dense calculations and fail in dimensions beyond $10^{4}$. We note that $P_{1}$ is the projection onto the probability simplex, also known as SparseMAX, for which more efficient, non-QP-based methods exist [Martins & Astudillo, 2016]. Results for $P_{2}$, representing projection onto “chains" with bounded links, exhibit similar scalability and are detailed in Appendix E.1.2.
@@
We further demonstrate the scalability of our approach on a bi-level optimization problem introduced in Kovalsky et al. [2020]:
@@
In this problem $v\in\mathbb{R}^{n\times 2}$ represents the $n$ vertex coordinates of a triangular mesh, $M\in\mathbb{R}^{n\times n}$ is a parameterized Laplacian, and $B$, $u$ and $C$ encode boundary conditions. $\mu^{*}(M)$ is the dual variable corresponding to linear inequalities of the lower-level problem. The results described in Kovalsky et al. [2020] imply that if $\mu^{*}(M)$ vanishes then $v^{*}(M^{*})$, at the optimal Laplacian $M^{*}$, represents a straight-edge intersection-free drawing of the mesh Tutte [1963]. The solution to (10) is visualized in Figure 5(a) for an example large-scale ant mesh. Additionally, Figure 5(b) shows that dQP scales more favorably with problem size compared to OptNet, QPLayer and SCQPTH; in particular, only dQP supports problems with over $10^{4}$ vertices.
@@ Appendix A Existing Methods for Differentiable Optimization
Name Programs Solvers Features Limitations OptNet [Amos & Kolter, 2017a] QP qpth GPU batchable. Re-uses factorization from forward for fast backward solve. Solver specific. Sparsity not supported. QPLayer [Bambade et al., 2024] QP ProxQP Supports infeasible QPs. Tightly couples to ProxQP. Solver specific. Sparsity not supported. SCQPTH [Butler, 2023] QP OSQP Differentiates ADMM updates. Supports infeasibility detection and automatic scaling. Solver specific. Sparsity not supported. Diffcp [Agrawal et al., 2019b] CP SCS, ECOS, Clarabel Supports sparse cone programs (CP). Supports specific conic solvers. CVXPYLayers [Agrawal et al., 2019a] CP diffcp Flexible problem formulation and solver choice. Supports sparsity. Inherits diffcp limitations. BPQP [Pan et al., 2024] QP,SOCP OSQP Computes derivatives by solving a second equality-constrained QP. Published implementation is incomplete. Alt-Diff [Sun et al., 2022] QP Custom ADMM Differentiates ADMM updates. Supports inexact solutions. Reportedly slow [Bambade et al., 2024]. LQP [Butler & Kwon, 2023] QP Custom ADMM Box constraints only. LPGD [Paulus et al., 2024] CP diffcp Modifies diffcp to handle degenerate derivatives. Inherits diffcp limitations. CVXPYgen [Schaller & Boyd, 2025] QP, SOCP OSQP Accelarates CVXPYlayers by C compilation of problem formulation. Solver specific. JAXopt-general [Blondel et al., 2022] User supplied User supplied Differentiates arbitrary implicit functions. User must manually provide KKT.. JAXopt-OSQP [Blondel et al., 2022] QP Custom GPU OSQP GPU batchable. Solver specific. JAXopt-CVXPYQP [Blondel et al., 2022] QP CVXPY Flexible problem formulation and solver choice. Inherits CVXPY limitations. Sparsity not supported. SDPRLayers [Holmes et al., 2024] CP CVXPY Flexible problem formulation and solver choice. Inherits CVXPY limitations. Torchopt [Ren* et al., 2023] User supplied User supplied Differentiates arbitrary implicit functions. User must manually provide KKT. Optax [DeepMind et al., 2020] Specialized Specialized Supports specialized problems/solvers, e.g., projections. DiffOpt.jl [Besançon et al., 2024] Various JuMP-supported solvers Dunning et al. [2017] Supports a variety of convex and non-convex problems. JuMP provides QP support only for COSMO, OSQP, and Ipopt Theseus [Pineda et al., 2022] NLS CHOLMOD, cudaLU, BaSpaCho Supports non-linear least squares (NLS). GPU batchable. Supports sparsity. Lacks hard constraints. OptNet-Sparse [Amos & Kolter, 2017c] QP qpth Published implementation is incomplete. OptNet-CVXPY [Amos & Kolter, 2017b] QP CVXPY Published implementation is incomplete. OSQPTh [Stellato & O’Donoghue, 2023] QP OSQP Published implementation is incomplete. DFWLayer [Liu et al., 2024] QP Frank-Wolfe Automatically differentiates an unrolled Frank-Wolfe solver. Solver specific. CombOptNet [Paulus et al., 2021] ILP Gurobi Differentiation of combinatorial solvers Blackbox-Backprop [Pogančić et al., 2020] MIP Gurobi MIP, Blossom V, Dijkstra Differentiation of combinatorial solvers

### v4

@@ 3.3 Extracting Derivatives from a QP Solver’s Solution
Figure 2: Illustration of active set differentiation. Left: a QP is shown by its quadratic level sets and polyhedral feasible set; the solution lies on a facet of the boundary; perturbations of the constraints lead to perturbations in the solution. Right: the perturbation of the solution remains the same when inactive constraints are eliminated.
@@
Locally Equivalent Equality-Constrained QP.
@@
A proof of this Theorem, based on the Basic Sensitivity Theorem [46], is provided in Appendix B, along with a calculation of the derivatives using differential matrix calculus [91, 78]. We note that this result is closely related to analyses studied in multi-parametric programming [21, 93, 106, 12, 82].
@@
Notably, for quadratic programming, the Basic Sensitivity Theorem allows us to bypass implicit differentiation techniques [70]. We emphasize that this does not imply that the general solution or its active set admit a closed-form expression. Rather, while we perform explicit differentiation, the implicit function theorem remains key in establishing the local equivalence between the original and reduced problems -— the resulting derivatives are the same, though the derivations differ. The derivatives in (7) are obtained via ordinary (explicit) differentiation of the closed-form solution to the reduced QP (6), whereas those in (4) arise from implicit differentiation of the full nonlinear KKT conditions (2), followed by pruning inactive constraints. This distinction, formalized in Theorem 3.1, underscores a critical computational insight: once a black-box solver returns the primal solution, the active set can be identified and the derivatives computed directly via (7). Moreover, if the solver returns only the primal solution, the corresponding dual variables can be recovered using (6). Since both (6) and (7) rely on the same KKT matrix $K_{J}$, a single matrix factorization (e.g., via SuperLU [74]) suffices for both steps, adding negligible overhead. These insights culminate in the core algorithm of dQP, summarized in Algorithm 1.
@@
Our approach enables compact and efficient gradient computation. The linear system in (7), used to compute both derivatives and dual variables, is symmetric and of reduced size. In contrast, implicit differentiation of the full KKT conditions (1) yields a significantly larger, asymmetric system (3). Beyond simplifying the derivative computation, our approach enables the use of fast, specialized linear solvers that exploit the reduced systems symmetric indefinite KKT matrix structure (e.g., using an LDL factorization as in QDLDL [108, 37]).
Figure 3: (a) Active-set parameter space, coloring regions where the active set is constant. (b) Condition number of the full and reduced linear systems along a line in parameter space.
Empirically, we observe that the reduced linear system (7) is often significantly better conditioned than its full counterpart. Figure 3 illustrates this with an example of a QP governed by two parameters $\theta=(\theta_{1},\theta_{2})$ of [106], calculated using DAQP [13]. Eliminating inactive constraints improves conditioning significantly.
Figure 3 also highlights a key challenge: computing derivatives near non-differentiable singularities where the active set changes and some constraints become weakly active. In such regions, implicit differentiation becomes severely ill-conditioned. Our method is also affected, primarily through difficulty in determining the correct active set at an approximate solution. Several strategies have been proposed for improving active set identification [33, 84, 28]. Our implementation includes an optional refinement heuristic (Appendix 6), though in all our experiments (Section 4), simple hard thresholding of the primal residual $r_{j}=(Cz^{*}-d)_{j}\geq-\epsilon_{J}$ proved robust.
@@
Our open-source implementation is available at https://github.com/cwmagoon/dQP. We implement dQP (Algorithm 1) as a fully differentiable PyTorch module [88], providing an intuitive interface for integrating differentiable QPs into machine learning and bi-level optimization workflows. The implementation supports both dense and sparse problems end-to-end, with appropriate QP and linear solver support. As a PyTorch module, (7) must be rendered as a backward pass; this is described in Appendix C. To ensure modularity, the forward pass supports any QP solver interfaced via the open-source qpsolvers library [32], which provides lightweight access to over 15 commercial and free open-source solvers and easily supports the integration of new ones. We likewise offer flexibility in the choice of linear solver for differentiation, including support for large-scale sparse solvers like Pardiso [103] and symmetric indefinite solvers like QDLDL [108, 37]. For users unsure of which solver to use, dQP includes a profiling tool to help identify the best-performing QP solver for a given problem. Additional implementation details such as constraint normalization, handling non-differentiability, and warm-starting options for bi-level optimization are discussed in Appendix D.
Table 1: Performance of differentiable QP methods for 129 sparse problems from the Maros-Meszaros (MM) dataset [79].
@@ 4 Experimental Results
We have extensively tested dQP to ensure its robustness and evaluate its performance against existing methods for differentiable quadratic programming. Notably, we demonstrate dQP’s strengths in handling large-scale structured and sparse problems, thus complementing custom differentiable GPU-batched solvers such as OptNet [5], which are optimized for solving many small, dense problems simultaneously. Given this focus, and considering the limited availability of state-of-the-art GPU-batchable QP solvers, we conduct our experiments on CPUs, similar to prior works [17, 29, 109, 1]. Our evaluation includes a large benchmark of over 2,000 challenging sparse and dense QPs taken from public and randomly generated problem datasets, designed to test dQPs robustness and performance. Additional experimental details are provided in Appendix E.
@@
We tested dQP on the QP Benchmark suite [31], focusing on 129 large-scale sparse problems from the standard Maros-Meszaros (MM) dataset [79], which are widely used as stress tests for QP solvers. We compared dQP against other differentiable QP methods available in the PyTorch framework: OptNet [5], QPLayer [17], SCQPTH [29], and CVXPYLayers [1]; using the authors’ open-source implementations, each paired with its respective solver. We did not include other general frameworks that either lack direct support for generic QPs in PyTorch, showed subpar performance in our preliminary tests (consistent with findings reported in prior work [17]), or rely on forward passes built on diffcp/CVXPYLayers (see Table 4). For dQP’s forward pass, we selected the top-performing QP solver for each problem, as identified by QP Benchmark [31], to demonstrate the importance of enabling differentiation for the solver that is best suited to each task.
Figure 4: Accuracy versus total forward/backward solve for the Maros-Meszaros dataset [79]. Each point represents a solved problem; point size illustrates dimension; problems solved solely by dQP are circled. The legend shows percentages of success rates; the solvers dQP used and their counts are PIQP 100, Gurobi 9, ProxQP 9, Clarabel 7, OSQP 2, MOSEK 1, QPALM 1.
The scatter plot in Figure 4 shows total runtime (forward + backward), duality gap (accuracy), and problem dimension (illustrated by point size) for each differentiable solver and problem. Aggregate performance across the full dataset and the subset of problems solvable by all methods is summarized in Table 1. CVXPYLayers failed on all but 17 small-scale problems and is therefore excluded from the reported statistics. The MM dataset proved especially challenging, with OptNet and SCQPTH solving fewer than 50% of the problems. In contrast, dQP solved all MM problems and was the only method to succeed on 38 of them (circled in the figure). Moreover, dQP achieved the best total forward/backward runtime and accuracy in 80% and 78% of all problems, respectively. It performed particularly well on large-scale instances (dimension over 1000), being fastest and most accurate in 97% and 93% of such cases. Additional experiments on 450 random dense QPs and 625 sparse QPs (dimensions $10$ to $10^{4}$) are provided in Appendix E.1.3, showing similarly strong performance.
@@
Results for $P_{1}$ are shown in Figure 1, demonstrating dQP’s scalability compared to OptNet and QPLayer. Other methods fail on all but small problems (see Appendix E.1.1). In dimensions greater than 2000, dQP outperforms competing methods by 2-3 orders of magnitude in both speed and accuracy. Competing methods are limited to dense calculations and fail in dimensions beyond $10^{4}$. We note that $P_{1}$ is the projection onto the probability simplex, also known as SparseMAX, for which more efficient, non-QP-based methods exist [80]. Results for $P_{2}$, representing projection onto “chains" with bounded links, exhibit similar scalability and are detailed in Appendix E.1.2.
@@
We further demonstrate the scalability of our approach on a bi-level optimization problem introduced in [69]:
@@
In this problem $v\in\mathbb{R}^{n\times 2}$ represents the $n$ vertex coordinates of a triangular mesh, $M\in\mathbb{R}^{n\times n}$ is a parameterized Laplacian, and $B$, $u$ and $C$ encode boundary conditions. $\mu^{*}(M)$ is the dual variable corresponding to linear inequalities of the lower-level problem. The results described in [69] imply that if $\mu^{*}(M)$ vanishes then $v^{*}(M^{*})$, at the optimal Laplacian $M^{*}$, represents a straight-edge intersection-free drawing of the mesh [112]. The solution to (10) is visualized in Figure 5(a) for an example large-scale ant mesh. Additionally, Figure 5(b) shows that dQP scales more favorably with problem size compared to OptNet, QPLayer and SCQPTH; in particular, only dQP supports problems with over $10^{4}$ vertices.
@@ 5 Conclusion
Acknowledgments
The authors gratefully acknowledge partial support from NSF grant DMS-2152289, FRG: Collaborative Research: Mathematical and Statistical Analysis of Compressible Data on Compressive Networks; the NSERC Discovery Grant RGPIN-2024-04605, Practical Neural Geometry Processing; and the FRQNT Établissement de la relève professorale grant 365040, Calcul rapide et léger des déformations à l’aide de réseaux neuronaux.
@@ Appendix A Existing Methods for Differentiable Optimization
Name Programs Solvers Features Limitations OptNet [5] QP qpth GPU batchable. Re-uses factorization from forward for fast backward solve. Solver specific. Sparsity not supported. QPLayer [17] QP ProxQP Supports infeasible QPs. Tightly couples to ProxQP. Solver specific. Sparsity not supported. SCQPTH [29] QP OSQP Differentiates ADMM updates. Supports infeasibility detection and automatic scaling. Solver specific. Sparsity not supported. Diffcp [2] CP SCS, ECOS, Clarabel Supports sparse cone programs (CP). Supports specific conic solvers. CVXPYLayers [1] CP diffcp Flexible problem formulation and solver choice. Supports sparsity. Inherits diffcp limitations. BPQP [86] QP,SOCP OSQP Computes derivatives by solving a second equality-constrained QP. Published implementation is incomplete. Alt-Diff [109] QP Custom ADMM Differentiates ADMM updates. Supports inexact solutions. Reportedly slow [17]. LQP [30] QP Custom ADMM Differentiates ADMM updates. Box constraints only. LPGD [90] CP diffcp Modifies diffcp to handle degenerate derivatives. Inherits diffcp limitations. CVXPYgen [102] QP, SOCP OSQP Accelarates CVXPYlayers by C compilation of problem formulation. Solver specific. JAXopt-general [24] User supplied User supplied Differentiates arbitrary implicit functions. User must manually provide KKT.. JAXopt-OSQP [24] QP Custom GPU OSQP GPU batchable. Solver specific. JAXopt-CVXPYQP [24] QP CVXPY Flexible problem formulation and solver choice. Inherits CVXPY limitations. Sparsity not supported. SDPRLayers [62] CP CVXPY Flexible problem formulation and solver choice. Inherits CVXPY limitations. Torchopt [96] User supplied User supplied Differentiates arbitrary implicit functions. User must manually provide KKT. Optax [39] Specialized Specialized Supports specialized problems/solvers, e.g., projections. DiffOpt.jl [22] Various JuMP-supported solvers [42] Supports a variety of convex and non-convex problems. JuMP provides QP support only for COSMO, OSQP, and Ipopt Theseus [92] NLS CHOLMOD, cudaLU, BaSpaCho Supports non-linear least squares (NLS). GPU batchable. Supports sparsity. Lacks hard constraints. OptNet-Sparse [7] QP qpth Published implementation is incomplete. OptNet-CVXPY [6] QP CVXPY Published implementation is incomplete. OSQPTh [107] QP OSQP Published implementation is incomplete. DFWLayer [77] QP Frank-Wolfe Automatically differentiates an unrolled Frank-Wolfe solver. Solver specific. CombOptNet [89] ILP Gurobi Differentiation of combinatorial solvers Blackbox-Backprop [94] MIP Gurobi MIP, Blossom V, Dijkstra Differentiation of combinatorial solvers